OpenStudy (anonymous):

Ron bought a piece of plastic that was 3 1/2 yd by 2 2/3 yd. What was the area of Ron's plastic? Express your answer in simplest form. A.21/16 yd2 B.6 1/6 yd2 C.6 1/3 yd2 D.9 1/3 yd2

OpenStudy (jojokiw3):

Multiply base times height. so (b)h=a thus. \[3\frac{ 1 }{ 2 }(2\frac{ 2 }{ 3 })\]

OpenStudy (anonymous):

or BxH=A

OpenStudy (anonymous):

3 1/2 x 2 2/3 to make it easier to understand

OpenStudy (anonymous):

So D

OpenStudy (jojokiw3):

Correct!

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

3 1/2 x 2 2/3 = 9 1/3